ANALYSIS AND DEPERTURBATION OF THE A^{2}Π and B^{2}Σ^{+} STATES OF CaF

\underline{LEONID A. KALEDIN}^{a}, Department of Chemistry, Emory University, Atlanta, GA 30322, U.S.A.; JONATHAN C. BLOCH, Lincoln Laboratory, Massachusetts Institute of Technology, Lexington, MA 02173, U.S.A.; MICHAEL C. McCARTHY, Center for Astrophysics, Harvard University, Cambridge, MA 02138, U.S.A.; ROBERT W. FIELD, Department of Chemistry, Massachusetts Institute of Technology, Cambridge, MA 02139, U.S.A.

A laser excitation spectrum of the (0,0) and (1,1) bands of the CaF A^{2}Π-X^{2}Σ^{+} system has been recorded. The present set of molecular constants for both A^{2}Π and B^{2}Σ^{+} states represent a vast improvement in precision over previously reported values^{b} because of larger number of lines position measurements and the use of a more reasonable effective Hamiltonian model which accounts explicitly for the A - B interaction via off-diagonal matrix elements. Accurate spectroscopic parameters for the X, A, and B states will be presented.

aSupport by AFOSR under grant F19628-93-C-0186 is gratefully acknowledged

bJ. Nakagawa, P. J. Domaille, T. C. Steimle, and D. O. Harris, J. Mol. Spectrosc. 70, 374-385 (1978); M. Dulick, P. F. Bernath, and R. W. Field, Canad. J. Phys. 58,703-712 (1980); J. Verges, C. Effantin, A. Bernard, A. Topouzkhanian, A. R. Allouche, J. d’Incan, and R. F. Barrow, J. Phys. B 26, 279-284 (1993)